OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 1) and (1, 10)?

OpenStudy (across):

Do you know the slope formula?\[m=\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

y=ax+b

OpenStudy (anonymous):

so a = m

OpenStudy (anonymous):

y= \[\frac{10-1}{1-(-2)}\]

OpenStudy (anonymous):

so \[m\]=3

OpenStudy (anonymous):

y=3x+b

OpenStudy (anonymous):

so now we sub x and y to get b

OpenStudy (across):

@moha_10, try writing everything in one post to make things ten times easier to read!

OpenStudy (anonymous):

okay i will

OpenStudy (anonymous):

to sub x by -2 and y by 10 so we got 10=3*(-2)+b so finally b= 16 so y = 3x+16